# Reassessing Sub-Neptune Structure, Radii, and Thermal Evolution

### v1

@@ 2.7 Silicate Mantle
$c_{\rm{Pm}}$ 1200 $\rm{J\ kg^{-1}\ K^{-1}}$ Specific heat capacity of silicate 29, 30 1
$c_{\rm{Pc}}$ 840 $\rm{J\ kg^{-1}\ K^{-1}}$ Specific heat capacity of Fe 24 2
$l_{\rm{Si}}$ 732.2 $\rm{kJ\ kg^{-1}}$ Specific latent heat of silicate 29, 30 3
$l_{\rm{Fe}}$ 1200 $\rm{kJ\ kg^{-1}}$ Specific latent heat of Fe 24 4
$k_{\rm{um}}$ 5 $\rm{W\ m^{-1}\ K^{-1}}$ Thermal conductivity of upper mantle 32 1
$k_{\rm{lm}}$ 10 $\rm{W\ m^{-1}\ K^{-1}}$ Thermal conductivity of lower mantle 32 1
$k_{\rm{m}}$ 4 $\rm{W\ m^{-1}\ K^{-1}}$ Thermal conductivity of magma 32 5
$k_{\rm{c}}$ 40 $\rm{W\ m^{-1}\ K^{-1}}$ Thermal conductivity of core 46 6
$\kappa_{\rm{m}}$ $\rm{1\times 10^{-6}}$ $\rm{m^{2}\ s^{-2}}$ Thermal diffusivity of mantle 32 7
$\eta_{\rm{m}}$ 100 $\rm{Pa\ s}$ Viscosity of magma 39 8
$Ra_{\rm{c}}$ 600 - Critical Rayleigh number 33 7
$\eta_{0}$ $1.9\times 10^{21}$ $\rm{Pa\ s}$ Reference viscosity for solid mantle 35 9
$E_{a}$ 162 $\rm{kJ\ mol^{-1}}$ Activation energy 36 9
$V_{a0}$ 1.4 $\rm{cm^{3}\ mol^{-1}}$ Reference activation volume 37 9
$T_{0}$ 1600 $\rm{K}$ Reference temperature 35 9
$p_{\rm{decay}}$ 1610 $\rm{GPa}$ pressure decay 37 9
$\tau_{\rm{CAI}}$ 3 Myr Time after CAI 31 10
$\phi_{\rm{crit}}$ 0.5 - critical melt fraction 38
$\Delta\phi$ 0.1 - transition width 38
@@
$\sigma$ $1\times{10^{6}}$ $\rm{S\ m^{-1}}$ Electrical conductivity of core 44 12
@@ 3.1 Contraction Due to Thermal Evolution
In the left panels of Figure 3, we show the evolution tracks of envelope specific entropy as a function of iron/rock mass, H/He mass fraction and incident bolometric flux. Our findings indicate that a sub-Neptune consistently contracts over time during post-boil-off evolution due to a decline in envelope specific entropy. We find the planets with either a more massive rock/iron core, more substantial envelope mass or a higher stellar insolation possess a warmer interior at a given age. This occurs because they retain a larger interior energy reservoir, even though their cooling luminosity is slightly higher. The difference is found to be small between planets with cold surfaces (see the $1F_{\oplus}$ and $10F_{\oplus}$ models in the bottom left).
@@
In most cases, the thermal evolution of a sub-Neptune’s envelope is primarily controlled by its own energy budget (as in Eq. 3) as well as the heat capacity of the rock/iron core, making it relatively insensitive to the solidification processes occurring deep within the rock/iron interior. However, exceptions arise for planets with both a substantial rock/iron interior($\geq 8M_{\oplus}$) and a small envelope mass (H/He mass fraction $\leq$ 0.2%). In such cases, the planet’s energy reservoir is predominantly governed by the rock/iron core, causing the envelope’s contraction rate to be directly influenced by the cooling rate of the mantle. As the bottom of the mantle solidifies, the mantle’s cooling rate accelerates, due to the reduced heat input from the iron core. This leads to a sudden increase in the envelope contraction rate. Figure 4 illustrates these shifts in contraction rates, which coincide with the phase transition at the CMB, as shown in the right panel. These model planets have a H/He mass fraction of 0.2%.
@@
In the right panels of Figure 3, we illustrate the evolution of the RCB pressure over time. As a planet cools, the RCB penetrates to deeper levels, indicating that the radiative atmosphere becomes a larger portion of the gaseous H/He layer. This behavior is anticipated, as the approximately constant RCB temperature ($\sim T_{\rm{eq}}$) must correspond to a higher pressure level of a colder adiabat. With an increase in incident bolometric flux, the disparity between the interior heat transport (which establishes the envelope adiabat) and the stellar heating (which determines the radiative atmospheric temperature) widens, causing the RCB to shift inward to depths exceeding several kilobars (bottom right). Additionally, planets with more massive rock/iron cores or a more massive envelope generally exhibit a shallower RCB (top and middle panels on the right) due to their warmer interiors. Among these influences, stellar heating exerts the most significant effect on the RCB location.
Note that the contribution of the radiative atmospheric mass has a significant evolutionary influence on the most highly irradiated planets (see Section 2.3) but not for the others. In the case of the $1000F_{\oplus}$ model (yellow), the radiative zone occupies a substantial portion of the H/He mass at older ages due to the deep RCB (bottom right). Consequently, the reduced envelope mass correspond to a smaller gravitational energy reservoir for cooling, leading to a faster contraction rate compared to the planets with negligible radiative atmospheric mass (those experiencing lower stellar insolation). This effect is particularly pronounced in planets with lower rock/iron mass and/or lower H/He mass.
@@ 3.3 Post-Boil-Off Mass Fraction
In this section, we present the mass fractions at the end of boil-off, which determine the maximum initial mass fraction permitted for subsequent evolution. Tables 5 and 6 show the mass fraction grids as functions of rock/iron mass and bolometric flux for pure H/He envelopes and envelopes with 50$\times$ solar metallicity, respectively. For low-mass and highly irradiated planets (bottom left), the envelopes are completely stripped due to boil-off mass loss, denoted by zeroes. Note that they may still hold a radiative atmosphere with a negligible H/He mass. Conversely, more massive and cooler planets (top right) retain all of their initial envelope mass post-boil-off and, therefore, could potentially start the later evolution with any amount of envelope mass.
@@ 3.4 Evolution of the Core and Mantle
The thermal states of the core and mantle determine their physical states. As cooling progresses, the interior temperature decreases, leading to solidification. The internal physical state of the silicate mantle, in turn, influences energy transport (cooling) rates, resulting in a coupling between these processes. We find that the silicate mantle solidifies in a bottom-up order for low-mass planets ($\leq 3M_{\oplus}$), where the local melt fraction monotonically increases with radius. However, for intermediate and high-mass planets, the melt fraction drops below 1 (indicating a transition out of the pure solid state) starting from the middle of the lower mantle rather than from the bottom. Consequently, the crystallization region quickly expands and propagates both upward and downward, leading to a large portion of the lower mantle entering a partially molten phase and solidifying nearly simultaneously. Notably, the density gradient remains negative regardless of the reversed gradient of the melt fraction at the mantle bottom, ensuring that the solid/liquid mixture is gravitationally stable. We observe that complete solidification always occurs first at the bottom rather than at the surface.
@@
For all modeled planets, we find that iron solidification begins from the center, gradually forming a solid inner core over time. The solidification onset timescale varies greatly among planets, ranging from a few Myr to a few Gyr. In the left panels, we show the fractional solid core radius normalized by the iron core radius. As expected, planets with either a higher rock/iron mass or more substantial envelope mass begin solidification at a later age due to their hotter mantles (which results from the hotter envelopes) that keep the iron core warm. An exception is observed for the $1M_{\oplus}$ planet (top), as explained below. None of the modeled planets’ cores are completely solidified by 10 Gyr. For planets with low and intermediate rock/iron mass (top), their fractional solid core radius tends to converge after a few Gyr, after which the solidification rate (the slope) becomes similar. In contrast, planets with varying envelope masses exhibit divergent solidification rates at later ages, with those having more massive envelopes solidifying at a faster pace, resulting in a wide dispersion in fractional solid core radii.
We argue that the physical state of the mantle significantly impacts the iron solidification timescale. If the bottom of the mantle remains in the liquid phase, iron solidification proceeds quickly due to vigorous mantle convection that extracts heat from the core. Once the bottom of the mantle becomes solid, the solidification rate sharply decreases, leading to changes in the slopes of the solidification curves. It is important to note that the solidification of the $1M_{\oplus}$ planet (top left) is delayed, indicating a warmer core condition compared to the $2-4M_{\oplus}$ planets, despite having a colder mantle (middle). This is because its mantle solidifies at a very young age from the bottom, slowing down the cooling of the iron core.
While our model effectively captures the relative solidification timescales among planets with different parameters, the quantitative details are expected to carry some uncertainty. We find that the iron solidification timescale is sensitive to the melting temperature, which is expected to be significantly influenced by impurities. In Figure 28, we explore various melting curves that represent fractions of the melting temperature for a pure composition to account for different levels of impurity. We find that the onset of iron solidification for a $5M_{\oplus}$ planet (solid line) ranges from 10 Myr to a few 100 Myrs, given a 30% difference in melting temperature. For a lower-mass planet (dashed, 1 $M_{\oplus}$), the difference in solidification onset timescale is even more pronounced, varying from just 1 Myr to over 1 Gyr. This large dispersion arises from early solidification at the bottom of the mantle, which greatly slows down core cooling, resulting in this sensitive behavior.
@@ 3.5.1 Traditional Treatments
However, the deep radiative atmosphere characterizes the location where the visible stellar radiation is mostly absorbed by the planet (Guillot, 2010), thus having a higher temperature $T_{\rm{deep}}$ than $T_{\rm{eq}}$. In contrast, at a higher altitude, the optically thin region is determined by a colder “skin temperature”. This can lead to uncertainty in radii calculations if a simplified T-P profile is used. Additionally, previous work often assumed local gravity to be constant, typically evaluated at the RCB, to further simplify computations. This neglect of variable gravity effects can result in an overestimation of local gravity at higher altitudes and, consequently, an underestimation of the planetary radius. Lastly, while the atmosphere is dominated by molecular hydrogen throughout the whole lower atmosphere as assumed in the previous models, a molecular-atomic transition occurring above the $\mu$bar level must be considered. This transition accounts for the decline in mean molecular weight; without considering it, the nanobar radius may be underestimated.
@@ 3.5.2 Boil-Off-Constrained Variable Gravity Effect in an Isothermal Atmosphere
Figure 9: We show the mass-radius diagrams for planets with a range of planetary age (top), H/He mass fraction (middle) and incident bolometric flux (bottom). Our default model has a H/He mass fraction of 1%, incident bolometric flux of $100F_{\oplus}$ and age of 1 Gyr. We compare models with variable gravity (solid) and constant gravity (dotted), both having an isothermal atmosphere. The rock/iron core radius is shown in dashed. For planets that are still boiling off, we mark the area with shaded color. For the post-boil-off evolution, the difference between the two models is small.
Figure 10: The atmospheric escape parameter $\lambda$ (top panel) and Mach number (bottom panel) as a function of planetary mass and H/He mass fraction. Both are evaluated at a 20 mbar pressure level and an age of 10 Myr. We compute the relative difference in the optical radius between the variable gravity model and constant gravity model (same models in Figure 9), with regions in parameter space marked by color-coded contours, each associated with a specific tolerance value. Note that these models do not include a boil-off phase. The gray shaded area corresponds to boil-off conditions, which should lose a substantial amount of H/He mass if mass loss is switched on. Thus they are excluded from the discussion.
@@ 3.5.3 Enhanced Variable Gravity Effect in a Non-Isothermal Atmosphere
In Figure 11, we show the local gravity (top), scale height (middle, normalized by $R_{\rm{RCB}}$) and radius (bottom) as a function of pressure (radius increases at a lower pressure in the radiative atmosphere), for a range of different planetary masses. The constant gravity models (dashed, same models as in Figure 9) are included in the bottom panel for comparison. All models have the same H/He mass fraction of 1%, which is allowed by a boil-off phase, an incident flux of 10 $F_{\oplus}$ and an age of 10 Myr. Planets with an inflated interior are found to be sensitive to the variable gravity effect, due to both the low gravity and large scale height above the RCB. Thus planets with low rock/iron mass, high stellar irradiation and/or high H/He envelope mass fraction are especially susceptible. Among planets constrained by boil-off, low-mass planets are more susceptible than high-mass planets, although the latter hold thicker envelopes. We find the secular thermal evolution of planets makes them less inflated, leading to a less important variable gravity effect. This effect, however, is modest compared to the influence from the interior composition and stellar heating. Moreover, the 1 nbar radius is usually a few scale heights above the optical radius, which consequently further reduces the gravity leading to a magnified variable gravity effect (bottom).
@@
Figure 12: Mass-radius diagrams that account for the atmospheric T-P profile and boil-off. In the top panel, we show isocompositional curves, comparing planets with different incident fluxes with those exposed to $10F_{\oplus}$ in red-purple and $100F_{\oplus}$ in blue (both having 1%). For the $10F_{\oplus}$ models, the H/He mass fractions are varied, ranging from 0.1% to 10%, with lighter colors indicating lower fractions and darker colors indicating higher ones. Planetary ages are represented by solid for younger planets (100 Myr) and dotted for older planets (10 Gyr). The curves are truncated due to boil-off constraints, indicating that such planets would lose their H/He envelopes and not retain the parameters specified. In the bottom panel, the mass-radius relationship is shown for a fixed incident flux, with the H/He mass fractions determined by a boil-off phase (see the initially H/He-rich configuration in Section 2.5 and Table 5). It is important to note that planets with small envelopes exhibit minimal shrinkage over 10 Gyr of thermal contraction due to the deep RCB effect, which results in a thicker radiative atmosphere. This behavior is discussed further in the text.

### v2

@@ 2.7 Silicate Mantle
$c_{\rm{Pm}}$ 1200 $\rm{J\ kg^{-1}\ K^{-1}}$ Specific heat capacity of silicate 31, 32 1
$c_{\rm{Pc}}$ 840 $\rm{J\ kg^{-1}\ K^{-1}}$ Specific heat capacity of Fe 26 2
$l_{\rm{Si}}$ 732.2 $\rm{kJ\ kg^{-1}}$ Specific latent heat of silicate 31, 32 3
$l_{\rm{Fe}}$ 1200 $\rm{kJ\ kg^{-1}}$ Specific latent heat of Fe 26 4
$k_{\rm{um}}$ 5 $\rm{W\ m^{-1}\ K^{-1}}$ Thermal conductivity of upper mantle 34 1
$k_{\rm{lm}}$ 10 $\rm{W\ m^{-1}\ K^{-1}}$ Thermal conductivity of lower mantle 34 1
$k_{\rm{m}}$ 4 $\rm{W\ m^{-1}\ K^{-1}}$ Thermal conductivity of magma 34 5
$k_{\rm{c}}$ 40 $\rm{W\ m^{-1}\ K^{-1}}$ Thermal conductivity of core 48 6
$\kappa_{\rm{m}}$ $\rm{1\times 10^{-6}}$ $\rm{m^{2}\ s^{-2}}$ Thermal diffusivity of mantle 34 7
$\eta_{\rm{m}}$ 100 $\rm{Pa\ s}$ Viscosity of magma 41 8
$Ra_{\rm{c}}$ 600 - Critical Rayleigh number 35 7
$\eta_{0}$ $1.9\times 10^{21}$ $\rm{Pa\ s}$ Reference viscosity for solid mantle 37 9
$E_{a}$ 162 $\rm{kJ\ mol^{-1}}$ Activation energy 38 9
$V_{a0}$ 1.4 $\rm{cm^{3}\ mol^{-1}}$ Reference activation volume 39 9
$T_{0}$ 1600 $\rm{K}$ Reference temperature 37 9
$p_{\rm{decay}}$ 1610 $\rm{GPa}$ pressure decay 39 9
$\tau_{\rm{CAI}}$ 3 Myr Time after CAI 33 10
$\phi_{\rm{crit}}$ 0.5 - critical melt fraction 40
$\Delta\phi$ 0.1 - transition width 40
@@
$\sigma$ $1\times{10^{6}}$ $\rm{S\ m^{-1}}$ Electrical conductivity of core 46 12
@@ 3.1 Contraction Due to Thermal Evolution
In the left panels of Figure 4, we show the evolution tracks of envelope specific entropy as a function of iron/rock mass, H/He mass fraction and incident bolometric flux. Our findings indicate that a sub-Neptune consistently contracts over time outside the boil-off phase, due to a decline in envelope specific entropy. We find the planets with either a more massive rock/iron core, more substantial envelope mass or a higher stellar insolation possess a warmer interior at a given age. This occurs because they retain a larger interior energy reservoir, even though their cooling luminosity is slightly higher. The difference is found to be small between planets with cold surfaces (see the $1F_{\oplus}$ and $10F_{\oplus}$ models in the bottom left).
@@
In most cases, the thermal evolution of a sub-Neptune’s envelope is primarily controlled by its own energy budget (as in Eq. 3), making it relatively insensitive to the solidification processes occurring deep within the rock/iron interior. However, exceptions arise for planets with both a substantial rock/iron interior($\geq 8M_{\oplus}$) and a small envelope mass (H/He mass fraction $\leq$ 0.2%). In such cases, the planet’s energy reservoir is predominantly governed by the rock/iron core, causing the envelope’s contraction rate to be directly influenced by both the cooling rate and the physical state of the mantle. As the bottom of the mantle solidifies, the mantle’s cooling rate accelerates, due to the reduced heat input from the iron core. This leads to a sudden increase in the envelope contraction rate. Figure 5 illustrates these shifts in both the envelope’s contraction rate (left panel) and the mantle’s cooling rate (right panel), which coincide with the phase transition at the CMB (circles). After the phase transition, the iron core begins to decouple from the thermal evolution of the mantle and envelope. These model planets have a H/He mass fraction of 0.2%.
@@
In the right panels of Figure 4, we illustrate the evolution of the RCB pressure over time. We examine planets with 1 to 20 $M_{\oplus}$, H/He mass fractions from 0.1% to 20%, and incident bolometric flux from 1 to 1000 $F_{\oplus}$. As a planet cools, the RCB penetrates to deeper levels, indicating that the radiative atmosphere becomes a larger portion of the gaseous H/He layer. This behavior is anticipated, as the approximately constant RCB temperature ($\sim T_{\rm{eq}}$) must correspond to a higher pressure level of a colder adiabat. With an increase in incident bolometric flux, the disparity between the interior heat transport (which establishes the envelope adiabat) and the stellar heating (which determines the radiative atmospheric temperature) widens, causing the RCB to shift inward to depths exceeding several kilobars (bottom right). Additionally, planets with more massive rock/iron cores or a more massive envelope generally exhibit a shallower RCB (top and middle panels on the right) due to their warmer interiors. Among the parameters chosen and within the ranges considered in this study, incident bolometric flux exerts the most significant effect on the RCB location.
Our results indicate that in extreme cases (bolometric flux $\gtrsim 100F_{\oplus}$, H/He mass fraction $\lesssim 0.2\%$, and core mass $\lesssim 3M_{\oplus}$), the incident bolometric radiation penetrates to the MEB after 1 Gyr, leading to an envelope-free, thick-atmosphere super-Earth, with the radiative atmosphere contributing more than 40% of its total radius.
Note that the contribution of the radiative atmospheric mass has a significant evolutionary influence on the most highly irradiated planets (see Section 2.3) but not for the others. In the case of the $1000F_{\oplus}$ model (yellow), the radiative zone occupies a substantial portion of the H/He mass at older ages due to the deep RCB (bottom right). Consequently, the reduced envelope mass corresponds to a smaller energy reservoir for cooling, leading to a faster contraction rate compared to the planets with negligible radiative atmospheric mass (those experiencing lower stellar insolation). This behavior is evident in the yellow curve in the bottom left panel, where the contraction rate accelerates significantly after a few Gyrs. This effect is particularly pronounced in planets with lower rock/iron mass and/or lower H/He mass.
@@ 3.3 Post-Boil-Off Mass Fraction
In this section, we present the mass fractions at the end of boil-off, which determine the maximum initial mass fraction permitted for subsequent evolution. Tables 5 and 6 show the mass fraction grids as functions of rock/iron mass and bolometric flux for pure H/He envelopes and envelopes with 50$\times$ solar metallicity, respectively. For low-mass and highly irradiated planets (bottom left), the envelopes are completely stripped due to boil-off mass loss, denoted by zeroes. Note that they may still hold a radiative atmosphere with a negligible H/He mass. Conversely, more massive and cooler planets (top right) retain all of their initial envelope mass post-boil-off and, therefore, could potentially start the later evolution with any amount of envelope mass. Constraining the initial envelope mass for these planets requires a planet formation model, which is beyond the scope of this study.
@@ 3.4 Evolution of the Core and Mantle
The thermal states of the core and mantle determine their physical states. As cooling progresses, the interior temperature decreases, leading to solidification. The internal physical state of the silicate mantle, in turn, influences energy transport (cooling) rates, resulting in a coupling between these processes. We find that the silicate mantle solidifies in a bottom-up order for low-mass planets ($\leq 3M_{\oplus}$), where the local melt fraction monotonically increases with radius. This behavior is illustrated in the top panel of Figure 30, which shows the melt fraction as a function of planetary radius in the mantle.
However, for intermediate and high-mass planets (bottom panel), the melt fraction drops below 1 (indicating a transition out of the pure liquid state) starting from the middle of the lower mantle rather than from the bottom. Consequently, the crystallization region quickly expands and propagates both upward and downward, leading to a large portion of the lower mantle entering a partially molten phase and solidifying nearly simultaneously. Notably, the density gradient remains negative regardless of the reversed gradient of the melt fraction at the mantle bottom, ensuring that the solid/liquid mixture is gravitationally stable. We observe that complete solidification always occurs first at the bottom rather than at the surface.
A middle-out solidification process could theoretically form a solid layer that inhibits cooling at the mantle bottom due to high viscosity, potentially leading to a persistent basal magma ocean (Lherm et al., 2023). However, the intermediate- and high-mass planets modeled in this study are likely not in this case, as nearly synchronized solidification appears to occur in the lower mantle once solid convection begins ($\phi<0.5$). Since our model does not explicitly capture energy transport within individual mantle mass shells, this behavior requires further investigation in future studies.
@@
For all modeled planets, we find that iron solidification begins from the center, gradually forming a solid inner core over time. The solidification onset timescale varies greatly among planets, ranging from a few Myr to a few Gyr. In the left panels, we show the fractional solid core radius normalized by the iron core radius. As expected, planets with either a higher rock/iron mass or more substantial envelope mass begin solidification at a later age due to their hotter mantles (which results from the hotter envelopes) that keep the iron core warm. An exception is observed for the $1M_{\oplus}$ planet (dashed curve with the latest onset of iron solidification in the top left panel), as explained below. None of the modeled planets’ cores are completely solidified by 10 Gyr. For planets with low and intermediate rock/iron mass (top), their fractional solid core radius tends to converge after a few Gyr, after which the solidification rate (the slope) becomes similar. In contrast, planets with varying envelope masses exhibit divergent solidification rates at later ages, with those having more massive envelopes solidifying at a faster pace, resulting in a wide dispersion in fractional solid core radii.
We argue that the physical state of the mantle significantly impacts the iron solidification timescale. If the bottom of the mantle remains in the liquid phase, iron solidification proceeds quickly due to vigorous mantle convection that extracts heat from the core. Once the bottom of the mantle becomes solid, the solidification rate sharply decreases, leading to changes in the slopes of the solidification curves. It is important to note that the solidification of the $1M_{\oplus}$ planet (top left) is delayed, indicating a warmer core condition compared to the $2-4M_{\oplus}$ planets, despite having a colder mantle (middle). This is because its mantle solidifies at a very young age from the bottom, slowing down the cooling of the iron core. Consequently, a positive feedback is triggered: the mantle experiences a reduced heat flux at the CMB due to the iron’s latent heat, which accelerates the mantle solidification process while delaying the solidification of the iron core. This effect becomes more pronounced in planets with less substantial envelopes and higher iron melting temperatures (see below) and may extend to more massive planets.
@@
For a comparison between our results, which incorporates the improved rock/iron core modeling, and previous models that assume isothermal cores, see Section 3.6.
[…]
Figure 9: Local melt fraction as a function of planetary radius in the silicate mantle. From top to bottom, the curves represent increasing time since the onset of mantle solidification (in units of Gyr). The top panel depicts a low-mass planet ($1M_{\oplus}$, 5%) that solidifies in a bottom-up manner, while the bottom panel illustrates a more massive planet ($4M_{\oplus}$, 1%) that begins solidification in the middle of the lower mantle. Both planets have stellar insolation of $100F_{\oplus}$. See text for further discussion.
@@ 3.5.1 Traditional Treatments
However, the deep radiative atmosphere characterizes the location where the visible stellar radiation is mostly absorbed by the planet (Guillot, 2010), thus having a higher temperature $T_{\rm{deep}}$ than $T_{\rm{eq}}$. In contrast, at a higher altitude, the optically thin region is determined by a colder “skin temperature” (see Figure 2). This can lead to uncertainty in radii calculations if a simplified T-P profile is used. Additionally, previous work often assumed local gravity to be constant, typically evaluated at the RCB, to further simplify computations. This neglect of variable gravity effects can result in an overestimation of local gravity at higher altitudes and, consequently, an underestimation of the planetary radius. Lastly, while the atmosphere is dominated by molecular hydrogen throughout the whole lower atmosphere as assumed in the previous models, a molecular-atomic transition occurring above the $\mu$bar level must be considered. This transition accounts for the decline in mean molecular weight; without considering it, the nanobar radius may be underestimated.
@@ 3.5.2 Boil-Off-Constrained Variable Gravity Effect in an Isothermal Atmosphere
Figure 11: We show the mass-radius diagrams for planets with a range of planetary age (top), H/He mass fraction (middle) and incident bolometric flux (bottom). Our default model has a H/He mass fraction of 1%, incident bolometric flux of $100F_{\oplus}$ and age of 1 Gyr. We compare models with variable gravity (solid) and constant gravity (dotted), both having an isothermal atmosphere. The rock/iron core radius is shown in dashed. For planets that are still boiling off, we mark the area with shaded color. At high rock/iron masses, the planetary radius is no longer constrained by boil-off (as discussed in Section 3.3), resulting in the truncation. In the bottom panel, note that the boil-off-prohibited region (shaded) extends to higher planetary masses with increasing flux. For the post-boil-off evolution, the difference between the two models is small.
Figure 12: The atmospheric escape parameter $\lambda$ (top panel) and Mach number (bottom panel) as a function of rock/iron mass and H/He mass fraction. Both are evaluated at a 20 mbar pressure level and an age of 10 Myr. We compute the relative difference in the optical radius between the variable gravity model and constant gravity model (same models in Figure 11), with regions in parameter space marked by color-coded contours, each associated with a specific tolerance value. Note that these models do not include a boil-off phase. The gray shaded area corresponds to boil-off conditions, which should lose a substantial amount of H/He mass if mass loss is switched on. Thus they are excluded from the discussion.
@@ 3.5.3 Enhanced Variable Gravity Effect in a Non-Isothermal Atmosphere
In Figure 13, we show the local gravity (top), scale height (middle, normalized by $R_{\rm{RCB}}$) and radius (bottom) as a function of pressure (radius increases at a lower pressure in the radiative atmosphere), for a range of different rock/iron core masses. The constant gravity models (dashed, same models as in Figure 11) are included in the bottom panel for comparison. All models have the same H/He mass fraction of 1%, which is allowed by a boil-off phase, an incident flux of 10 $F_{\oplus}$ and an age of 10 Myr. Planets with an inflated interior are found to be sensitive to the variable gravity effect, due to both the low gravity and large scale height above the RCB. Thus planets with low rock/iron mass, high stellar irradiation and/or high H/He envelope mass fraction are especially susceptible. Among planets constrained by boil-off, low-mass planets are more susceptible than high-mass planets, although the latter hold thicker envelopes. We find the secular thermal evolution of planets makes them less inflated, leading to a less important variable gravity effect. This effect, however, is modest compared to the influence from the interior composition and stellar heating. Moreover, the 1 nbar radius is usually a few scale heights above the optical radius, which consequently further reduces the gravity leading to a magnified variable gravity effect (bottom).
@@
Figure 14: Mass-radius diagrams that account for the atmospheric T-P profile and boil-off. In the top panel, we show isocompositional curves, comparing planets with different incident fluxes with those exposed to $10F_{\oplus}$ in red-purple and $100F_{\oplus}$ in blue (both having 1%). For the $10F_{\oplus}$ models, the H/He mass fractions are varied, ranging from 0.1% to 10%, with lighter colors indicating lower fractions and darker colors indicating higher ones. Planetary ages are represented by solid for younger planets (100 Myr) and dotted for older planets (10 Gyr). The curves are truncated due to boil-off constraints, indicating that such planets would lose their H/He envelopes and not retain the parameters specified. The prohibited region is shown in gray (see discussion in Section 3.3). In the bottom panel, the mass-radius relationship is shown for a fixed incident flux, with the H/He mass fractions determined by a boil-off phase (see the initially H/He-rich configuration in Section 2.5 and Table 5). It is important to note that planets with small envelopes exhibit minimal shrinkage over 10 Gyr of thermal contraction due to the deep RCB effect, which results in a thicker radiative atmosphere. This behavior is discussed further in the text.